OpenStudy (anonymous):

Create two sets of three triangle lengths: one set that does create a right triangle and another that does not. Show and explain why the first set does create a right triangle and why the second set does not. PLEASE HELP!!!!

OpenStudy (cwrw238):

right - angled triangle : lengths 3,4,5 non right-angled: lengths 7,8 , 10 the first set creates a rt angled triangle because 5^2 = 3^2 + 4^2 (pythagoras theorem) in the second triangle 10^2 not = 7^2 + 8^2 so not right-angled

OpenStudy (cwrw238):

any more help needed?

OpenStudy (anonymous):

no but thank you :)